OpenStudy (anonymous):

Graph the circle with the center at (3,-2) which also passes through the point (0,2). Find the radius.

OpenStudy (campbell_st):

well the form of the circle you need is \[(x - h)^2 + (y - k)^2 = r^2\] (h, k) is the centre of the circle so your circle is \[(x - 3)^2 + (y + 2)^2 = r^2\] now substitute the point x = 0 and y = 2 to find r^2 then take the square root to find the radius

OpenStudy (anonymous):

thank you